OpenStudy (anonymous):

F(x)= -6x^6(x)^1/2 + -3/x^2(x)^1/2 find f'(x)

OpenStudy (anonymous):

Is this the function? \[ -6x^6(x)^{\frac{1}{2}} + \frac{-3}{x^2(x)^\frac{1}{2}}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[x^6x^\frac{1}{2} = x^{6+\frac{1}{2}} = x^\frac{13}{2}\] Differentiate this first.

OpenStudy (anonymous):

\[\frac{1}{x^2x^\frac{1}{2}}=\frac{1}{x^{2+\frac{1}{2}}}=\frac{1}{x^{\frac{5}{2}}}=x^{-\frac{5}{2}}\]And differentiate this.

OpenStudy (anonymous):

Use: \[\frac{d}{dx}x^n = nx^{n-1}\]